HUBBLE SPACE TELESCOPE DAILY REPORT #2117 PERIOD COVERED: 0000Z (UTC) 05/05/98 - 0000Z (UTC) 05/06/98 Daily Status Report as of 126/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 7277 (Calibration of the Cepheid P-L Relation with Observations of the Maser-Host Galaxy NGC) The WF/PC-2 was used to make observations of the maser-host galaxy NGC-4258 to study its Cepheid variable stars. These data will help refine the calibration of the Cepheid period-luminosity relationship. The observations were completed as scheduled, and no problems were reported. 1.2 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.3 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed STIS/CCD/MA2 7393 (Black Hole A0620-00 and Advection-Dominated Accretion) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to make observations of soft X-ray transient V616-Mon. These data might provide the most clear-cut evidence presently available for the existence of black holes. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.7 Completed Four Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.8 Completed Three Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.9 Completed NIC/1/2/3 7854 (The Onset of Axisymmetry in Proto-Planetary Nebulae) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the axissysmitry of the proto-planetary nebula CIT6. The high-resolution imaging and multicolor polarimetric capabilities of NICMOS comprise a superb combination for studying the processes by which evolved (approximately one solar mass) stars shed their outer layers to form planetary nebulae. The observations were completed as planned, and no problems were reported. 1.10 Completed STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.11 Completed WF/PC-2 7409 (Resolving sdB Binary Systems) The WF/PC-2 was used to make observations to try and resolve the sdB class binary system Star-13114. This is a continuation of the approved Cycle 6 Proposal No. 6351, which due to unlucky circumstances did not succeed in obtaining any data until 8/30/96. The observations were completed as scheduled, and no anomalies were reported. 1.12 Completed STIS/CCD 7666 (CCD Linearity and Shutter Stability) The Space Telescope Imaging Spectrograph (CCD) was used to make some CCD linearity and shutter stability calibration checks. CCDs can suffer from non-linearities at low, high {or both} intensity. Such effects can be subtle, but important contributor to error budgets for routine science observations and are sometimes recognized only after limitations are detected in careful analyses of science data. The observations were completed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 7386 (A Snapshot Survey of Nova Shells) The WF/PC-2 was used to make observational images of a super nova shell in order to detect and measure its expanding gaseous shell. These observations will trace the distribution of hot and cool gas in the nova shells, and show their degree of inhomogeneity on both large and small scales. The observations were completed as scheduled, and no problems were noted. 1.14 Completed NIC/1/2/3 7829 (Mapping H_2 Emission around T Tauri Stars) The Near Infrared Camera (Cameras One, Two, and Three) was to be used to study the distribution of circumstellar dust and molecular {H_2} gas around six classical T Tauri stars {cTTS}, probing both the near circumstellar environment {the accretion disk} as well as the environment further out where strong winds from these stars may be interacting with remnants of the cloud material from which they formed. The primary goal is to map the distribution of the molecular gas around cTTS. As described in HSTAR 6459 and in 2.1, the acquisition and the two-reacquisitions for this approximate three-and-one-half-hour proposal failed and the take data flag remained down. 1.15 Completed STIS/CCD 7911 (STIS Galactic Plane Pure Parallel Program) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude survey observations in parallel with the prime science observations. Short exposures with the 50CCD imaging mode will provide the wavelength zero-point calibration. These data will be placed immediately into the Hubble Data Archive. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed NIC/1/2/3 7896 (A NIR Snapshot Survey of Ultraluminous IR Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make snapshot survey observations of the ultraluminous infrared galaxy QDOTULIR06. Nearly all of these galaxies show evidence for a recent tidal interaction, and we will search for evidence of the merger progenitors near the center of each galaxy, a task made significantly easier with NIR imaging. The observations were completed as planned, and no anomalies were noted. 1.17 Completed WF/PC-2 7470 (A Snapshot Survey of Galactic Globular Clusters) The WF/PC-2 was used to make snap shot survey observations of the crowded center of the galactic globular cluster NGC-6273. The resulting color-magnitude diagram will include all stars down to about one magnitude below the main-sequence turnoff. The observations were completed as scheduled, and no problems were reported. 1.18 Completed STIS/MA1 7604 (Cycle 7 MAMA Dark Measurements) 30|35\; The Space Telescope Imaging Spectrograph (MA1) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.19 Completed WF/PC-2/STIS/CCD/NIC/1/2/3 7307 (The Formation and Evolution of Rich Star Clusters in the LMC) The WF/PC-2, the Space Telescope Imaging Spectrograph (CCD), and the Near Infrared Camera (Cameras One, Two and Three) were used to make observations of the rich star clusters which are located within the Large Magellanic Cloud in order to study their formation and evolution. These observations will help us understand the origin and evolution of rich star clusters in our own galaxy. The observations were completed as scheduled, and no anomalies were reported. 1.20 Completed S/C 7954 (FGS-1R Optimization) Desorption and gravity release effects have modified the morphology and modulation of the FGS1R Transfer Function. This test will accomplish a re-optimization of the performance of FGS1R by adjusting the FF3 mirror until the S Curve at the center of the field of view is comparable to the optimized S Curve obtained during SMOV. This test requires special data downlinks and uplinks for real-time commanding of the FF3. The test concluded with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 12 Successful : 11 Scheduled Reacquisitions : 8 Successful : 6 The acquisition at 125/1533Z and the re-acquisitions at 125/1639Z and 125/1709Z failed when the search radius limit was exceeded. HSTAR 6459 was written. The proposal described in 1.14 was affected. 2.2 FHST Updates: Scheduled : 30 Successful : 30 2.3 Operations Notes: The SSR EDAC counts were cleared using ROP SR-1. As a result of the acquisition failures described above, an ARU/PRT was performed at 125/1805Z. The following acquisitions were successful. Also, using ROP NS-11, the NICMOS error counter was cleared. The optimization of FGS-1R, as described in 1.20, continues. Today, an Internal Test Source (ITS) acquisition was performed at 125/2330Z. The Articulated Mirror Assembly (AMA) was commanded eight steps in the reverse direction at 125/2354Z, as directed by operations request, with no anomalous behavior reported. Another ITS acquisition was executed at 126/0203Z, along with two POS exposures. Two S-curves were produced. After an analysis meeting today, another AMA move will be performed. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH